# HyWA: Hypernetwork Weight Adapting Personalized Voice Activity Detection

### v1

@@ Abstract
Personalized Voice Activity Detection (PVAD) systems activate only in response to a specific target speaker by incorporating speaker embeddings from enrollment utterances. Unlike existing methods that require architectural changes, such as FiLM layers, our approach employs a hypernetwork to modify the weights of a few selected layers within a standard voice activity detection (VAD) model. This enables speaker conditioning without changing the VAD architecture, allowing the same VAD model to adapt to different speakers by updating only a small subset of the layers. We propose HyWA-PVAD, a hypernetwork weight adaptation method, and evaluate it against multiple baseline conditioning techniques. Our comparison shows consistent improvements in PVAD performance. HyWA also offers practical advantages for deployment by preserving the core VAD architecture. Our new approach improves the current conditioning techniques in two ways: i) increases the mean average precision, ii) simplifies deployment by reusing the same VAD architecture.
Index Terms—  Personalized Voice Activity Detection, Hypernetwork, Weight Adaptation, Speaker Conditioning.
@@ 1 Introduction
Voice Activity Detection (VAD) is typically the first module in many speech processing pipelines, serving as a gating mechanism to activate downstream components, such as Automatic Speech Recognition [1]. VAD’s role becomes especially critical on edge device, where minimizing power consumption is the key concern. VAD helps conserve computational and energy resources by ensuring that subsequent modules are activated only during speech segments. Conventional VAD systems operate at the 10–20ms frame level, making binary {“speech”, “non-speech”} decisions for each incoming audio frame. These systems are expected to be lightweight, fast, and robust across a wide range of acoustic environments to be viable for real-world use, especially on resource-constrained devices [2, 3].
@@
Many edge devices are designed for single-user scenarios, where voice-based functionalities, such as voice assistants, are intended to be accessed only by the device owner. Typically, this is achieved by combining VAD with speaker verification. However, such multistage systems are computationally inefficient and introduce latency, particularly when employing large speaker verification models that require processing longer audio segments [4].
@@
There is growing interest in personalized VAD (PVAD) systems [4, 5] to enable more seamless and efficient interaction. PVAD models are trained to activate only in response to a specific user’s voice. Operation of PVAD models typically includes an enrollment phase, during which the user records a few samples of their own voice. These samples are then used to compute speaker embeddings and representations, which are subsequently fed into the PVAD model.
@@
Some of the most successful PVAD models in the literature modify traditional VAD models to incorporate speaker identity directly into the VAD processing pipeline [6]. These approaches typically rely on a speaker embedding extracted from a short enrollment utterance of the target speaker using a pre-trained speaker encoder. This embedding is then injected into the VAD model in various ways, such as feature concatenation at intermediate layers, bias modulation, or activation scaling. The process of injecting the speaker information into a VAD model is commonly referred to as speaker conditioning [7]. One of the most widely used speaker conditioning mechanisms is the Feature-wise Linear Modulation (FiLM) layer [8]. The FiLM layer conditions the VAD model on speaker embedding by applying scale and shift affine transformations on intermediate feature maps. FiLM-based speaker conditioning effectively personalizes the model without requiring an explicit speaker classification objective [5, 9, 10].
@@
A significant limitation of speaker conditioning mechanisms is the need to retrain the VAD model, or modify the architecture of the base VAD. Given the importance of VAD systems in production, such architectural changes and retraining are infeasible for deployment on edge devices. We propose a new speaker conditioning approach that leverages hypernetwork [11] to personalize an existing VAD. A hypernetwork is an auxiliary model that conditions the VAD on speaker information. Specifically, the hypernetwork generates weights for a small number of layers within the existing VAD model. This hypernetwork adapts to a target speaker while preserving the core VAD architecture.
@@
Hypernetworks have previously been used in various contexts, such as text-to-image generation [12], image editing [13], and meta learning [14]. However, hypernetworks are unexplored in the context of VAD personalization to the best of our knowledge. Our design offers a lightweight, modular, and practical solution for deploying PVAD capabilities on existing VAD models. This approach involves minimal overhead without compromising robustness or deployment constraints.
@@
We propose a novel speaker conditioning mechanism HyWA, that achieves the followings:
@@
  • provides a novel conditioning method to PVADs by employing a hypernetwork as a user-specific VAD parameter adapter.
  • improves the existing speaker conditioning methods for PVADs.
@@
  • benefits from a trained VAD model and does not require retraining from scratch.
Furthermore, we intend to release our full training and inference code pipeline to provide an open-source baseline for PVADs, comparing several speaker conditioning methods.
@@ 2 Methodology
Personalized voice activity detection (PVAD) requires integrating the target speaker’s enrollment data with the acoustic features of the input audio. A key challenge in PVAD development is how to use these inputs altogether so that the personalization is achieved without losing model performance. We address this issue by employing a hypernetwork.
@@ 2.1 Personalization
VADs are evolved to personalized versions by extending their input towards adding user-specific features. This extension is implemented by concatenating, adding, multiplying, or feature-wise linear modulation (FiLM) [15, 16]. This requires designing a new PVAD architecture which differs from the common VADs. Evolving VADs to PVADs through personalizing their weights is natural and hassle-free at the product development level. Instead of changing the input to obtain a PVAD, we employ the same VAD and modify the weights for that specific user through a hypernetwork. A hypernetwork, introduced in [11], is a metamodel that generates the weights of a primary model, using metadata. Instead of direct optimization of primary model parameters, the hypernetwork learns a mapping from metadata input to the parameter space of the primary model. In our case, we use the hypernetwork to generate the personalized parameters for the VAD model, using the enrollment voice of the user as the hypernetwork metadata.
Assume a VAD model is $\mathcal{M}_{\mathbf{w}}(\cdot)$ and the personalized VAD model, PVAD, for individual $k$ is $\mathcal{M}_{\mathbf{w}_{k}}(\cdot)$. We implicitly assume an individual $k$ has a speech characteristic vector $\mathbf{s}$, also called speaker embedding. Our hypernetwork focuses on generating specific weights for a given $\mathbf{s}$. We may reparameterize $\mathbf{w}_{k}\!=\!\mathbf{w}\!+\!\Delta\mathbf{w}_{k},$ and instead of learning $\mathbf{w}_{k}$, learn $\Delta\mathbf{w}_{k}$. This type of reparameterization is well studied in the context of parameter-efficient fine-tuning, also called adapters [17]. Adapters allow switching between PVAD and VAD by setting $\Delta\mathbf{w}_{k}\!=\!\mathbf{0}$. We propose to personalize only a subset of the VAD weights, i.e. enforcing a sparse structure on $\Delta\mathbf{w}$. This mechanism is particularly advantageous for personalization, because hypernetwork generates individual-specific weights by only modifying certain effective layers of the primary VAD model [18]. The training is performed on a set of individuals $i\!\in\!\{1,\ldots,n\}$ with speaker embedding $\mathbf{s}_{i}$, and audio features $\mathbf{a}_{i}$.
Fig. 1: An overview of the training pipeline for HyWA-PVAD. The hypernetwork $\mathcal{H}_{\theta}$ produces user-specific subset of VAD weights $\Delta\mathbf{w}$ based on speaker embedding $\mathbf{s}_{i}$. The VAD takes $\Delta\mathbf{w}$ along with audio features $\mathbf{a}$, and outputs PVAD labels $y$.
@@ 2.2 Training
The training of a common VAD model $y_{t}\!\sim\!\mathcal{M}_{\mathbf{w}}(\mathbf{a}_{t})$ includes optimization of a loss function over the audio signal $\mathbf{a}_{t}$ at a time stamp $t$, with a binary label $y_{t}$ that represents a human speech indicator {“speech”, “non-speech”}. In the sequel, we drop indices $t,k$ when there is no danger of confusion. PVAD training, however, is more complex. One may pre-train a VAD model first and then start personalization throughout our proposed process or simply train all parameters from scratch. Our PVAD training is decomposed into a general VAD block, parameterized in $\mathbf{w}$ and a subset of personalized weights $\Delta\mathbf{w}$.
The training pipeline is visualized in Figure 1. For individual $k$ at time stamp $t$, the PVAD model receives the audio features $\mathbf{a}$ and its subset parameters generated through the hypernetwork $\Delta\mathbf{w}\!\sim\!\mathcal{H}_{\theta}(\mathbf{s}).$ Note that the hypernetwork $\mathcal{H}_{\theta}(\cdot)$ only receives the speaker embedding $\mathbf{s}$. The whole training data include the speaker embedding $\mathbf{s}$, audio features $\mathbf{a}$ and the PVAD ternary labels $y$ that involve {“non-speech (ns)”, “target speaker speech (tss)”, “non-target speaker speech (ntss)”}. VAD model receives the user-specific VAD subset weights $\Delta\mathbf{w}$ from $\mathcal{H}_{\theta}$, the audio feature $\mathbf{a}$, and outputs the PVAD ternary labels $y$. A cross-entropy loss function over ternary labels, then, is used to train the whole set of parameters $(\theta,\mathbf{w},\Delta\mathbf{w})$ simultaneously.
@@ 2.3 Inference
    Deployment: feeding $\Delta\mathbf{w}$ into $\mathcal{M}$ to obtain a PVAD model $\mathcal{M}_{\mathbf{w}+\Delta\mathbf{w}}$, see Figure 2(a),
@@
    Usage: feeding audio features $\mathbf{a}$ at every time stamp into $\mathcal{M}_{\mathbf{w}+\Delta\mathbf{w}}$ to obtain personalized ternary labels $y$, see Figure 2(b).
@@
More precisely, after the training is performed, $(\theta,\mathbf{w})$ are learned. Evolving VAD towards PVAD for each individual requires feeding $\mathcal{M}$ with its user-specific weights $\Delta\mathbf{w}$, which is performed at the enrollment stage. The enrollment includes feeding the hypernetwork $\mathcal{H}_{\theta}(\mathbf{s})$ in which $\mathbf{s}$ is the speaker embedding for that user. This step generates the user-specific PVAD weights $\Delta\mathbf{w}$. Note that the enrollment needs to be performed only once per user to provide the VAD model with $\mathcal{M}_{\mathbf{w}+\Delta\mathbf{w}}(\cdot)$ to personalize, see Figure 2(a). This user-specific PVAD deployment is easy, because only the weights of the VAD model are modified with no architectural change, but other PVADs include new architectures and their deployment requires additional coding effort. In our proposed approach, returning to the regular VAD is easy by nulling the user-specific weights $\Delta\mathbf{w}\!=\!\mathbf{0}$, and combining {“tss”, “ntss”} into a single category {“speech”} during usage. The evolved PVAD model only requires the audio feature at each instance $\mathbf{a}$ to predict the PVAD ternary labels $y$, see Figure 2(b).
(a) Enrollment and Deployment
(b) Usage
Fig. 2: An overview of the inference pipeline for HyWA-PVAD; where (a) shows the enrollment and deployment stages, executed once per user through cloud-device communication, and (b) illustrates the usage, executed on the device after enrollment and deployment stages.
Fig. 3: Speaker conditioning methods for PVAD. Each approach illustrates a distinct strategy for integrating speaker information with acoustic features to enable personalization.
3 Experiments and Results
3.1 Dataset Construction
Multi-speaker datasets with natural speakers and associated speakers’ identity information are scarce [4]. To overcome this limitation, we construct a simulated multi-speaker dataset following the methodology proposed in [10]. Specifically, we uniformly sample 1 to 3 utterances from individual speakers, randomly selecting one utterance to represent the target speaker. These utterances are concatenated to form multi-speaker segments, simulating real conversational scenarios.
We use the “train-clean-100” subset of the LibriSpeech dataset [19]. Speech transcripts provided by LibriSpeech are employed to generate labels via forced alignment [20]. Framewise speaker labels are derived from the speaker identity metadata included in this dataset. To enhance model robustness, we employ multistyle training (MTR) [21] by augmenting the training data with noise from the MUSAN dataset [22]. The noise is augmented at signal-to-noise ratio (SNR) levels ranging from -5 to 20 dB, with 5 dB increments. During training, we augment only with MUSAN’s free-sound subset. At test time, seen noise uses free-sound, while unseen noise uses MUSAN’s sound-bible subset, which is excluded from training. We incorporate room acoustics using recorded room impulse responses (RIRs) as described in [23]. Model performance is validated on LibriSpeech “dev-clean” subset during training and evaluated on “test-clean” subset after training. Training continues until no significant improvement is observed on the validation set. Individuals in the training, validation, and test sets are non-overlapping to ensure fair evaluation.
3.2 Model Architecture
Our VAD model $\mathcal{M}$ is inspired by [15] that include a 2-layer LSTM with 64 hidden units. We add a two-layer perceptron before the LSTM block, and a single-layer perceptron after the LSTM block to boost the VAD personalization capacity. This gives a VAD architecture with $\approx 85k$ parameters on-device. The hypernetwork $\mathcal{H}$ receives the speaker embedding as input, and gives the user-specific parameters $\Delta\mathbf{w}$ to personalize the VAD model $\mathcal{M}$. We chose a 4-layer perceptron with GeLU activations, normalization, and a skip connection at each layer, totaling $\approx 3.6M$ parameters in the cloud.
3.3 Evaluation Metrics
Following [4, 10], we evaluate the performance of the model by computing the average precision score (AP) for each class, with the mean average precision (mAP) serving as the primary evaluation metric. The mAP score is determined by taking the average of the AP scores across all classes.
@@ 3.3 Evaluation Metrics
Non-Speech (ns) Target Speaker Speech (tss) Non-Target Speaker Speech (ntss)

### v2

# Ghazvini Nejad Jafarzadeh Asl Edraki Sadeghi Yu Asgharian Partovi Nia

@@ Abstract
Personalized Voice Activity Detection (PVAD) systems activate only in response to a specific target speaker. Speaker-conditioning methods are employed to inject information about the target speaker into a VAD pipeline, to achieve personalization. Existing speaker-conditioning methods typically modify the inputs or activations of a VAD model. We propose an alternative perspective to speaker conditioning. Our approach, HyWA, employs a hypernetwork to generate personalized weights for a few selected layers of a standard VAD model. We evaluate HyWA against multiple baseline speaker-conditioning techniques using a fixed backbone VAD. Our comparison shows consistent improvements in PVAD performance. This new approach improves the current speaker-conditioning techniques in two ways: i) increases the mean average precision, ii) facilitates deployment by reusing the same VAD architecture.
keywords:
Personalized Voice Activity Detection, Hypernetworks, Target-Speaker Detection, Speaker Conditioning
@@ 1 Introduction
Voice Activity Detection (VAD) is typically the first module in many speech processing pipelines, serving as a gating mechanism to activate downstream components, such as Automatic Speech Recognition [ramirez2007voice]. VAD's role is critical on edge devices, where minimizing power consumption is the key concern. VAD helps to conserve computational and energy resources by ensuring that subsequent modules are activated only during speech segments. Conventional VAD systems operate at the 10–20ms frame level, making binary {``speech", ``non-speech"} decisions for each incoming audio frame. These systems are expected to be lightweight, fast, and robust across a wide range of acoustic environments to be viable for real-world use [yadav2022hardware, OURS_MLSP2025_tinynoiserobustvoiceactivity].
@@
Many edge devices are designed for single-user scenarios, where voice-based functionalities, such as voice assistants, are intended to be accessed only by the device owner. Typically, this is achieved by combining VAD with speaker verification. However, such multistage systems are computationally inefficient and introduce latency, particularly when employing large speaker verification models that require processing longer audio segments [personal_vad_1].
@@
There is growing interest in personalized VAD (PVAD) systems [personal_vad_1, personal_vad_2] to enable more seamless and efficient interaction. PVAD models are trained to activate only in response to a specific user's voice. Operation of PVAD models typically includes an enrollment phase, during which the user records a few samples of their own voice. These samples are then used to compute speaker embeddings and representations, which are subsequently fed into the PVAD model.
@@
Some of the most successful PVAD models in the literature modify traditional VAD models to incorporate speaker identity directly into the VAD processing pipeline [medennikov2020target]. These approaches typically rely on a speaker embedding extracted from a short enrollment utterance of a target speaker using a pre-trained speaker encoder. This embedding is then injected into the VAD model in various ways, such as feature concatenation at intermediate layers, bias modulation, or activation scaling. The process of injecting the speaker information into a VAD model is commonly referred to as speaker conditioning [wang2018voicefilter]. One of the widely-used speaker conditioning mechanisms is the Feature-wise Linear Modulation (FiLM) layer [perez2018film]. The FiLM layer conditions the VAD model on speaker embedding by applying scale and shift affine transformations on intermediate feature maps. FiLM-based speaker conditioning effectively personalizes the model without requiring an explicit speaker classification objective [personal_vad_2, kang2023svvad, SSL-pvad-baseline].
@@
A significant limitation of speaker-conditioning mechanisms is the need to retrain the VAD model, or modify the architecture of the base VAD. Given the importance of VAD systems in production, such architectural changes and retraining are infeasible for deployment on edge devices. We propose a new perspective to speaker conditioning, that leverages hypernetwork [ha2017hypernetworks] to personalize an existing VAD. Our hypernetwork is an auxiliary model that conditions the VAD on speaker information. Specifically, the hypernetwork generates weights for a small number of layers within the existing VAD model. This hypernetwork adapts to a target speaker while preserving the core VAD architecture.
@@
Hypernetworks have been used previously in various contexts, such as text-to-image generation [ruiz2024hyperdreambooth], image editing [alaluf2022hyperstyle], and meta learning [zhao2020meta]. However, they are unexplored in the context of VAD personalization to the best of our knowledge. Our design offers a lightweight, modular, and practical solution for deploying PVAD capabilities on existing VAD models. This approach involves minimal overhead without compromising robustness or deployment constraints.
@@
We propose a novel speaker-conditioning mechanism, HyWA, that achieves the followings
@@
  • Provides a novel conditioning method for PVAD by employing a hypernetwork that generates user-specific VAD weights.
  • Outperforms the existing speaker conditioning methods for PVAD in mean average precision.
@@
Furthermore, upon acceptance we will release the complete training and inference pipeline, including code, configuration files, and evaluation scripts. This provides an open-source PVAD baseline for comparing speaker-conditioning methods lacking in the literature. Public release during review is restricted by internal policy related to intellectual property.
@@ 2 Methodology
Personalized voice activity detection (PVAD) requires integrating the target speaker’s enrollment data with the acoustic features of the input audio. A key challenge in PVAD development is how to use these inputs altogether so that the personalization is achieved without losing PVAD performance. We address this issue by employing a hypernetwork.
@@ 2.1 Personalization
VADs are evolved to personalized versions by extending their input towards adding user-specific features. This extension is implemented by concatenating, adding, multiplying, or feature-wise linear modulation (FiLM) [SSL-pvad-baseline-journalVersion, comparative-analysis-personalized-voice]. This requires designing a new PVAD architecture which differs from the common VADs. Evolving VADs to PVADs through personalizing their weights is natural and hassle-free at the product development level. Instead of changing the input to obtain a PVAD, we generate the VAD weights for that specific user through a hypernetwork. A hypernetwork, introduced in [ha2017hypernetworks], is a metamodel that generates the weights of a primary model, using metadata. Instead of direct optimization of primary model parameters, the hypernetwork learns a mapping from metadata input to the parameter space of the primary model. In our case, we use the hypernetwork to generate the personalized parameters for the VAD model, using the enrollment voice of the user as the hypernetwork metadata.
Assume a VAD model is $\mathcal{M}_{\mathbf{w}}(\cdot)$ and the personalized VAD model, PVAD, for individual $k$ is $\mathcal{M}_{\mathbf{w}_{k}}(\cdot)$. We implicitly assume an individual $k$ has a speech characteristic vector $\mathbf{s}_{k}$, also referred to as speaker embedding. Our hypernetwork aims to generate speaker-specific weights conditioned on $\mathbf{s}_{k}$. We reparameterize the personalized weights as $\mathbf{w}_{k}=\mathbf{w}+\Delta\mathbf{w}_{k}$, where $\Delta\mathbf{w}_{k}$ captures a speaker-specific residual update over the shared VAD parameter $\mathbf{w}$. In HyWA, $\Delta\mathbf{w}_{k}$ is generated in a single forward pass of a hypernetwork conditioned on $\mathbf{s}_{k}$, unlike parameter-efficient fine-tuning methods. This mechanism is particularly advantageous for personalization, because hypernetwork generates individual-specific weights by only modifying certain effective layers of the primary VAD model [chauhan2023hypernetworks]. Training is performed over a set of individuals $k\in\{1,\ldots,n\}$ with speaker embeddings $\mathbf{s}_{k}$ and audio features $\mathbf{a}_{k}$.
Figure 1: An overview of the training pipeline for HyWA-PVAD. The hypernetwork $\mathcal{H}_{\theta}$ produces user-specific subset of VAD weights $\Delta\mathbf{w}$ based on speaker embedding $\mathbf{s}_{i}$. The VAD takes $\Delta\mathbf{w}$ along with audio features $\mathbf{a}$, and outputs PVAD labels $y$.
@@ 2.2 Training
The training of a common VAD model $y_{t}\sim\mathcal{M}_{\mathbf{w}}(\mathbf{a}_{t})$ includes optimization of a loss function over the audio signal $\mathbf{a}_{t}$ at a time stamp $t$, with a binary label $y_{t}$ that represents a human speech indicator {``speech'', ``non-speech''}. In the sequel, we drop indices $t,k$ when there is no danger of confusion.
The training pipeline is visualized in Figure 1. In addition to hypernetwork parameters $\theta$, the general block in VAD parameterized as a subset of $\mathbf{w}$, can be trained from scratch, or pre-trained VAD weights can be used instead. The personalization process is applied throughout our proposed approach, where the the subset of personalized weights $\Delta\mathbf{w}$ are generated by hypernetwork and are directly applied to user-specific layers.
As shown in Figure 1, for individual $k$ at time stamp $t$, the PVAD model receives the audio features $\mathbf{a}$ and its subset parameters generated through the hypernetwork $\Delta\mathbf{w}\sim\mathcal{H}_{\theta}(\mathbf{s}).$ Note that the hypernetwork $\mathcal{H}_{\theta}(\cdot)$ only receives the speaker embedding $\mathbf{s}$. The whole training data include the speaker embedding $\mathbf{s}$, audio features $\mathbf{a}$ and the PVAD ternary labels $y$ that involve {``non-speech (ns)", ``target speaker speech (tss)", ``non-target speaker speech (ntss)"}. VAD model receives the user-specific VAD subset weights $\Delta\mathbf{w}$ from $\mathcal{H}_{\theta}$, the audio feature $\mathbf{a}$, and outputs the PVAD ternary labels $y$. A cross-entropy loss function over ternary labels, then, is used to train the whole set of parameters $(\theta,\mathbf{w},\Delta\mathbf{w})$ simultaneously.
(a) Enrollment and Deployment
(b) Usage
Figure 2: An overview of the inference pipeline for HyWA-PVAD; where (a) shows the enrollment and deployment stages, executed once per user through cloud-device communication, and (b) illustrates the usage, executed on the device after enrollment and deployment stages.
Figure 3: Speaker conditioning methods for PVAD. Each approach illustrates a distinct strategy for integrating speaker information with acoustic features to enable personalization.
@@ 2.3 Inference
    Deployment: feeding $\Delta\mathbf{w}$ into $\mathcal{M}$ to obtain a PVAD model $\mathcal{M}_{\mathbf{w}+\Delta\mathbf{w}}$, see Figure 2a,
@@
    Usage: feeding audio features $\mathbf{a}$ at every time stamp into $\mathcal{M}_{\mathbf{w}+\Delta\mathbf{w}}$ to obtain personalized ternary labels $y$, see Figure 2b.
Although HyWA does not aim to reduce latency, usage-time inference follows the same computational path as standard PVAD conditioning methods. Any extra computation is incurred only once offline during enrollment, when $\mathcal{H}_{\theta}(\mathbf{s})$ generates $\Delta\mathbf{w}$.
@@
More precisely, after the training is performed, $(\theta,\mathbf{w})$ are learned. Evolving VAD towards PVAD for each individual requires feeding $\mathcal{M}$ with its user-specific weights $\Delta\mathbf{w}$, which is performed at the enrollment stage. The enrollment includes feeding the hypernetwork $\mathcal{H}_{\theta}(\mathbf{s})$ in which $\mathbf{s}$ is the speaker embedding for that user. This step generates the user-specific PVAD weights $\Delta\mathbf{w}$. Note that the enrollment needs to be performed only once per user to provide the VAD model with $\mathcal{M}_{\mathbf{w}+\Delta\mathbf{w}}(\cdot)$ to personalize, see Figure 2a. This user-specific PVAD deployment is easy, because only the weights of the VAD model are modified with no architectural change, but other PVADs include new architectures and their deployment requires additional coding effort. In our proposed approach, returning to the regular VAD is easy by nulling the user-specific weights $\Delta\mathbf{w}=\mathbf{0}$, and combining {``tss", ``ntss"} into a single category {``speech"} during usage. The evolved PVAD model only requires the audio feature at each instance $\mathbf{a}$ to predict the PVAD ternary labels $y$, see Figure 2b.
@@
    Non-Speech (ns)     Target Speech (tss)     Non-Target Speech (ntss)
@@
3 Experiments and Results
3.1 Dataset Construction
There are a few benchmark datasets for PVADs [personal_vad_1]. To overcome this limitation, we construct a simulated multi-speaker dataset following the methodology proposed in [SSL-pvad-baseline]. Specifically, we uniformly sample 1 to 3 utterances from individual speakers, randomly selecting one utterance to represent the target speaker. These utterances are concatenated to form multi-speaker segments, simulating real conversational scenarios.
We use the ``train-clean-100'' subset of the LibriSpeech dataset [panayotov_Librispeech_2015]. Speech transcripts provided by LibriSpeech are employed to generate labels via forced alignment [mcauliffe2017montreal]. Framewise speaker labels are derived from the speaker identity metadata included in this dataset. To enhance model robustness, we employ multistyle training (MTR) [prabhavalkar2015automatic] by augmenting the training data with noise from the MUSAN dataset [musan2015]. The noise is augmented at signal-to-noise ratio (SNR) levels ranging from 5-5 to $20$ dB, with $5$ dB increments. During training, we augment only with MUSAN’s free-sound subset. At test time, seen noise uses free-sound, while unseen noise uses MUSAN’s sound-bible subset, which is excluded from training. We use the same SNR range and sampling procedure to create test sets. Moreover, we incorporate room acoustics using recorded room impulse responses (RIRs) as described in [ko2017study]. Model performance is validated on LibriSpeech ``dev-clean'' subset during training and evaluated on ``test-clean'' subset after training. Training continues until no significant improvement is observed on the validation set. Individuals in the training, validation, and test sets are non-overlapping to ensure fair evaluation.
3.2 Model Architecture
Our VAD model $\mathcal{M}$ is inspired by [SSL-pvad-baseline-journalVersion] and includes a 2-layer LSTM with 64 hidden units. We add a two-layer perceptron before the LSTM block and a single-layer perceptron after the LSTM block to boost personalization capacity, yielding an on-device VAD with $\approx 85$k parameters. For speaker conditioning, following [SSL-pvad-baseline] we extract 256-dimensional speaker embeddings using the Resemblyzer VoiceEncoder [icassp2018_resemblyzerpaper]. Each speaker embedding is computed from at least 3 seconds of enrollment audio in total, formed by concatenating one or more short utterances from that speaker. The hypernetwork $\mathcal{H}$ takes this speaker embedding as input and outputs user-specific parameters $\Delta\mathbf{w}$ that personalize $\mathcal{M}$. We use a 4-layer perceptron with GeLU activations, normalization, and a skip connection at each layer, totaling $\approx 3.6$M parameters in the cloud. We restrict personalization to linear layers to keep the design and analysis simple.
3.3 Evaluation Metrics
Following [personal_vad_1, SSL-pvad-baseline], we evaluate the performance of the model by computing the average precision score (AP) for each class, with the mean average precision (mAP) serving as the primary evaluation metric. The mAP score is determined by taking the average of the AP scores across all classes.